# Neutrino masses and mixed dark matter from doublet and singlet scalars

### v1

@@ 1 Introduction
There is no doubt that we need to go beyond the Standard Model due to nonzero masses for neutrinos. The scotogenic model Ma (2006) is among the mechanisms for generating neutrino masses, and it requires an inert scalar doublet (one more scalar doublet with a vanishing VEV) and three right-handed neutrinos. Unlike the high-scale seesaw mechanism Minkowski (1977); Yanagida (1979); Gell-Mann et al. (1979); Mohapatra and Senjanovic (1980), neutrino masses are generated by radiative corrections with the inert scalar doublet and the RH neutrinos running in loops. So, there is no need to take new particles in the scotogenic model to be much heavier than the weak scale, and the lighter neutral scalar field of the extra Higgs doublet is a good candidate for dark matter (DM). Then, there are testable signatures of the inert scalar doublet by dark matter direct detection, electroweak precision tests and collider experiments.
@@ 5 Benchmark models
In this section, we consider some benchmark models depending on the mixing angles between the DM scalars and the corresponding neutrino masses in our model.
@@
The dark matter candidate depends on the mass ordering for the $Z_{2}$-odd dark sector scalars. For $m_{H_{0}}>m_{a},m_{s}$, the lighter singlet-like scalar is a DM candidate: $A_{2}$ can be a DM candidate for $\mu>0$, for which $m_{a}<m_{s}$; $H_{2}$ can be a DM candidate for $\mu<0$, for which $m_{s}<m_{a}$; $H_{2}$ and $A_{2}$ are degenerate in mass for $\mu=0$, so the singlet-like complex scalar can be a DM candidate. On the other hand, for $m_{H_{0}}<m_{a},m_{s}$, either $H_{1}$ and $A_{1}$ can be a DM candidate.
@@ 5.1 Scenario I: no dark matter mixings
Figure 2: Scenario I: singlet CP-odd scalar dark matter. (a) Variation of the relic density as a function of $M_{DM}$. The color code indicates the variation of the $A_{2}A_{2}h_{2}$ coupling. (b) Branching fractions of DM annihilation cross sections as a function of $M_{DM}$. We chose a set of the parameters, $\lambda_{2}=\lambda_{S}=\lambda_{H_{1}\varphi}=\lambda_{H_{2}S}=0$, $\lambda_{3}=0.1,\lambda_{4}=-0.01,\lambda_{H_{2}\varphi}=\lambda_{\varphi}=1$, $\lambda_{H_{1}S}=[10^{-4},1]$, and $\lambda_{S\varphi}=[10^{-2},4\pi]$. The mass parameters are taken to $m_{2}^{2}=10^{9}$ GeV, $\mu=-1$ GeV, $m_{h_{2}}=150$ GeV and $m_{S}^{2}=[1,10^{8}]$ GeV.
For $\lambda_{S\varphi}\neq 0$ and/or $\mu\neq 0$, if open kinematically, the annihilation channel, $H_{2}H_{2}\to h_{2}h_{2}$ plays an important role to set the observed relic density. To explore this, we consider the parameters as $\mu=-1$ GeV and $m_{h_{2}}=150$ GeV, $\lambda_{H_{1}S}=[10^{-4},1]$ and $\lambda_{S\varphi}=[10^{-2},4\pi]$. The values of other model parameters are given in the second column of Table. 2. We show the relic density as a function of $M_{DM}$ for the varying effective coupling, $|\lambda_{H_{2}H_{2}h_{2}}|=(2\mu-\lambda_{S\varphi}v_{\varphi})$, in Fig. 2. We show the observed relic density value by the black line. The chosen parameter space satisfies the DD limit, EW precision data as well as theoretical constraints such as perturbative unitarity and vacuum stability. The charged Higgs scalar does not alter the $h\to\gamma\gamma$ rate as it is considered to be heavier than $10^{4}$ GeV.
@@
Figure 3: Scenario I: singlet CP-odd scalar dark matter. Parameter space satisfying $1\%-100\%$ of the observed relic density. The choice of the parameters is the same as in Fig. 2.
@@ 5.2 Scenario II: small dark matter mixings
If $\lambda_{H_{2}\varphi}=0$, the DM phenomenology is the same as in the inert doublet model. But, for $\lambda_{H_{2}\varphi}\neq 0$, the extra annihilation channel, $A_{2}A_{2}\to h_{2}h_{2}$, plays an important role to set the observed relic density, if open kinematically.
@@
In Fig. 5, we present the variation of $\Omega_{A_{1}}h^{2}$ in the $M_{DM}=m_{A_{1}}$ vs $|\lambda_{A_{1}A_{1}h_{1}}|$ plane, assuming $\lambda_{H_{2}\varphi}=0$. Here, we demanded $A_{1}$ to build up $1\%-100\%$ of the observed relic density while being consistent with the direct detection bound on $|\lambda_{A_{1}A_{1}h_{1}}|=|\lambda_{L}|v_{H}$ with $\lambda_{L}=\lambda_{3}+\lambda_{4}$. As evident from the plot, the DM relic density is overabundant for $M_{\rm DM}\gtrsim 10^{3}$ GeV and $|\lambda_{A_{1}A_{1}h_{1}}|<1$ GeV.
@@ 5.4 Scenario IV: large mass gap for singlet dark scalars
In Fig. 8, we show the relic density as a function of dark matter mass for a varying $\lambda_{A_{2}A_{2}h_{2}}$. Here, all the points are allowed by the direct detection constraints and other above mentioned constraints. In Fig. 8, we show variation of branching ratio for the process $DM\leavevmode\nobreak\ DM\to h_{2}h_{2}$ in $M_{DM}$ vs $\Omega_{A_{2}}h^{2}$ plane. In the parameter space with $\Omega h^{2}\leq 0.12$, the dark matter annihilation is dominated by $A_{2}A_{2}\to h_{2}h_{2}$, which makes up $>95\%$ contribution to the total annihilation cross section. The contract term contribution to $A_{2}A_{2}\to h_{2}h_{2}$ is suppressed for a small value of $\lambda_{S\varphi}$, but the s-channel and t-channel contributions to the corresponding DM annihilation cross section can be sizable due to a sizable value of the trilinear coupling, $|\lambda_{A_{2}A_{2}h_{2}}|=|2\mu-\lambda_{S\varphi}v_{\varphi}|$.

### v2

@@ 1 Introduction
There is no doubt that we need to go beyond the Standard Model due to nonzero masses for neutrinos. The scotogenic model Ma (2006) is among the mechanisms for generating neutrino masses, and it requires an inert scalar doublet (one more scalar doublet with a vanishing VEV) and three right-handed neutrinos. Unlike the high-scale seesaw mechanism Minkowski (1977); Yanagida (1979); Gell-Mann et al. (1979); Mohapatra and Senjanovic (1980); Schechter and Valle (1980), neutrino masses are generated by radiative corrections with the inert scalar doublet and the RH neutrinos running in loops. So, there is no need to take new particles in the scotogenic model to be much heavier than the weak scale, and the lighter neutral scalar field of the extra Higgs doublet is a good candidate for dark matter (DM). Then, there are testable signatures of the inert scalar doublet by dark matter direct detection, electroweak precision tests and collider experiments.
@@ 5 Benchmark models
In this section, we consider some benchmark models depending on the mixing angles between the DM scalars and the corresponding neutrino masses in our model. We generate the model files from SARAH-v4.15.2 Staub (2014) and use SPheno-v4.0.3 Porod (2003) for calculation of particle spectra and decays, and micrOmegas-v6.0.5 Alguero et al. (2024) for the calculation of dark matter observables.
@@
In the following discussion, we consider the case that the RH neutrinos are heavier than the lightest neutral scalar field in our model  [^2] See Ref. de la Vega et al. (2024) for the case with the RH neutrino dark matter.. Then, the candidate for scalar dark matter depends on the mass ordering for the $Z_{2}$-odd dark sector scalars. For $m_{H_{0}}>m_{a},m_{s}$, the lighter singlet-like scalar is a DM candidate: $A_{2}$ can be a DM candidate for $\mu>0$, for which $m_{a}<m_{s}$; $H_{2}$ can be a DM candidate for $\mu<0$, for which $m_{s}<m_{a}$; $H_{2}$ and $A_{2}$ are degenerate in mass for $\mu=0$, so the singlet-like complex scalar can be a DM candidate. On the other hand, for $m_{H_{0}}<m_{a},m_{s}$, either $H_{1}$ and $A_{1}$ can be a DM candidate.
@@ 5.1 Scenario I: no dark matter mixings
Figure 2: Scenario I: singlet CP-even scalar dark matter. (a) Variation of the relic density as a function of $M_{DM}$. The color code indicates the variation of the $A_{2}A_{2}h_{2}$ coupling. (b) Branching fractions of DM annihilation cross sections as a function of $M_{DM}$. We chose a set of the parameters, $\lambda_{2}=\lambda_{S}=\lambda_{H_{1}\varphi}=\lambda_{H_{2}S}=0$, $\lambda_{3}=0.1,\lambda_{4}=-0.01,\lambda_{H_{2}\varphi}=\lambda_{\varphi}=1$, $\lambda_{H_{1}S}=[10^{-4},1]$, and $\lambda_{S\varphi}=[10^{-2},4\pi]$. The mass parameters are taken to $m_{2}^{2}=10^{9}$ GeV, $\mu=-1$ GeV, $m_{h_{2}}=150$ GeV and $m_{S}^{2}=[1,10^{8}]$ GeV.
For $\lambda_{S\varphi}\neq 0$ and/or $\mu\neq 0$, if open kinematically, the annihilation channel, $H_{2}H_{2}\to h_{2}h_{2}$ plays an important role to set the observed relic density. To explore this, we consider the parameters as $\mu=-1$ GeV and $m_{h_{2}}=150$ GeV, $\lambda_{H_{1}S}=[10^{-4},1]$ and $\lambda_{S\varphi}=[10^{-2},4\pi]$. The values of other model parameters are given in the second column of Table. 2. We show the relic density as a function of $M_{DM}$ for the varying effective coupling, $|\lambda_{H_{2}H_{2}h_{2}}|=|2\mu-\lambda_{S\varphi}v_{\varphi}|$, in Fig. 2. We show the observed relic density value by the black line. The chosen parameter space satisfies the DD limit, EW precision data as well as theoretical constraints such as perturbative unitarity and vacuum stability. The charged Higgs scalar does not alter the $h\to\gamma\gamma$ rate as it is considered to be heavier than $10^{4}$ GeV.
@@
Figure 3: Scenario I: singlet CP-even scalar dark matter. Parameter space satisfying $1\%-100\%$ of the observed relic density. The choice of the parameters is the same as in Fig. 2.
@@ 5.2 Scenario II: small dark matter mixings
If $\lambda_{H_{2}\varphi}=0$, the DM phenomenology is the same as in the inert doublet model. But, for $\lambda_{H_{2}\varphi}\neq 0$, the extra annihilation channel, $A_{1}A_{1}\to h_{2}h_{2}$, plays an important role to set the observed relic density, if open kinematically.
@@
In Fig. 5, we present the variation of $\Omega_{A_{1}}h^{2}$ in the $M_{DM}=m_{A_{1}}$ vs $|\lambda_{A_{1}A_{1}h_{1}}|$ plane, assuming $\lambda_{H_{2}\varphi}=0$. Here, we demanded $A_{1}$ to build up $1\%-100\%$ of the observed relic density while being consistent with the direct detection bound on $|\lambda_{A_{1}A_{1}h_{1}}|=|\lambda_{L}|v_{H}$ with $\lambda_{L}=\lambda_{3}+\lambda_{4}$. As evident from the plot, the DM relic density is overabundant for $M_{\rm DM}\gtrsim 10^{3}$ GeV and $|\lambda_{A_{1}A_{1}h_{1}}|<1$ GeV. The reason is the following. For heavy DM masses $\gtrsim 1\,{\rm TeV}$, the DM annihilations into longitudinal components of W, Z bosons are dominant. According to the Goldstone equivalence theorem (for $\sqrt{s}\simeq 2m_{A_{1}}\gg m_{W,Z}$), $A_{1}A_{1}\to Z_{L}Z_{L}$ is proportional to $\lambda_{L}$ and $A_{1}A_{1}\to W_{L}W_{L}$ is proportional to $\lambda_{4}$ in our case. Since $\lambda_{L}$ is bounded by direct detection whereas $\lambda_{4}$ is bounded by electroweak precision data (as it leads to a mass splitting between $m_{H^{\pm}}$ and $m_{H_{0},A_{0}}$ within the inert doublet). So, both $A_{1}A_{1}\to Z_{L}Z_{L}$ and $A_{1}A_{1}\to W_{L}W_{L}$ are suppressed for DM mass greater than $1\,{\rm TeV}$.
@@ 5.4 Scenario IV: large mass gap for singlet dark scalars
In Fig. 8, we show the relic density as a function of dark matter mass for a varying $\lambda_{A_{2}A_{2}h_{2}}$. Here, all the points are allowed by the direct detection constraints and other above mentioned constraints. In Fig. 8, we show variation of branching ratio for the process $DM\leavevmode\nobreak\ DM\to h_{2}h_{2}$ in $M_{DM}$ vs $\Omega_{A_{2}}h^{2}$ plane. In the parameter space with $\Omega h^{2}\leq 0.12$, the dark matter annihilation is dominated by $A_{2}A_{2}\to h_{2}h_{2}$, which makes up $>95\%$ contribution to the total annihilation cross section. The contact term contribution to $A_{2}A_{2}\to h_{2}h_{2}$ is suppressed for a small value of $\lambda_{S\varphi}$, but the s-channel and t-channel contributions to the corresponding DM annihilation cross section can be sizable due to a sizable value of the trilinear coupling, $|\lambda_{A_{2}A_{2}h_{2}}|=|2\mu-\lambda_{S\varphi}v_{\varphi}|$. For instance, for the choice of parameters in Fig. 8, $\mu=100\,{\rm GeV}$, $v_{\varphi}=10^{4}\,{\rm GeV}$ and $\lambda_{S\varphi}=[10^{-4},1]$, we can take $\lambda_{A_{2}A_{2}h_{2}}$ to be as large as $10^{4}\,{\rm GeV}$, so the contact term contribution to $A_{2}A_{2}\to h_{2}h_{2}$ becomes subdominant for a small value of $\lambda_{S\varphi}$.